Burke Tires just paid a dividend of D0 = $2.00. Analysts expect the company's dividend to grow by 30% this year, by 10% in Year 2, and at a constant rate of 5% in Year 3 and thereafter. The required return on this low-risk stock is 9.00%. What is the best estimate of the stock's current market value?

Select the correct answer.

a. $67.98

b. $70.94

c. $69.46

d. $70.20

e. $68.72
Answer #1

Year Calculations 0 1 =$2*1.3 2=2.60*1.10 3 =2.86*1.05 Dividend $ 2.0000 $ 2.6000 $ 2.8600 $ 3.0030 Terminal Value at the end

Correct Option :a. $67.98